OpenStudy (handsomechickenlover):

I need help with this question! WILL MEDAL AND FAN! Add parentheses for this problem to equal 19. 5+4^2+6-2^2-1

OpenStudy (handsomechickenlover):

\[5+4^{2}+6-2^{2}-1\]

OpenStudy (handsomechickenlover):

HELP! MEDALS SHALL BE GIVEN

OpenStudy (danieldjpon3):

ok ill try

OpenStudy (danieldjpon3):

ummm

OpenStudy (danieldjpon3):

umm

OpenStudy (danieldjpon3):

i rlly need medals

OpenStudy (danieldjpon3):

lol im new

OpenStudy (danieldjpon3):

kk back to math

OpenStudy (danieldjpon3):

so what do i do again?

OpenStudy (danieldjpon3):

Meh i give up

OpenStudy (danieldjpon3):

lol

OpenStudy (handsomechickenlover):

okay.

OpenStudy (danieldjpon3):

do (5+(4x4)) + (6-(2x2) - 1 = 22 I think :P

OpenStudy (danieldjpon3):

lol

OpenStudy (danieldjpon3):

what grade is this

OpenStudy (danieldjpon3):

like high school

OpenStudy (handsomechickenlover):

This is 8th grade

OpenStudy (handsomechickenlover):

The question asked for the answer to be 19 though I can't seem to get it.

OpenStudy (danieldjpon3):

oh ok

OpenStudy (danieldjpon3):

im in 7th grade

OpenStudy (mathgeek289):

Come On. XD Just Answer It.

OpenStudy (mathgeek289):

Add parentheses for this problem to equal 19. 5+4^2+6-2^2-1 5+4^2+6−2^2−1

OpenStudy (mathgeek289):

We Need To Solve Them First

OpenStudy (danieldjpon3):

Yes

OpenStudy (mathgeek289):

5+4^2+6−2^2−1 4^2 = 16 2^2 = 4 5 + 16 + 6 - 4 - 1

OpenStudy (danieldjpon3):

that is 22

OpenStudy (danieldjpon3):

hmmm

OpenStudy (mathgeek289):

5 + 16 = 21 21 + 6 - 4 - 1 21 + 6 = 27. 27 - 4 - 1

OpenStudy (danieldjpon3):

22

OpenStudy (danieldjpon3):

ya

OpenStudy (mathgeek289):

27 - 4 = 23 23 - 1 = 22 So, We Need To Add Parenthesis To Make The First Expression Equal To 22.

OpenStudy (danieldjpon3):

kk well ima go and help someone else, bc i prob wont get medal so ye

OpenStudy (danieldjpon3):

ummm

OpenStudy (danieldjpon3):

if we do the equation we get 22

OpenStudy (danieldjpon3):

so mybe OMG I GOT IT !!!!

OpenStudy (danieldjpon3):

5+4^2+6-2^2-1

OpenStudy (danieldjpon3):

lets do dis

OpenStudy (danieldjpon3):

(2-1) ^2

OpenStudy (danieldjpon3):

is 2 right?

OpenStudy (danieldjpon3):

So if we do

OpenStudy (johnweldon1993):

\[\large (2-1)^2 \cancel{=} 2\]

OpenStudy (danieldjpon3):

???

OpenStudy (johnweldon1993):

\[\large (2-1)^2 = (1)^2 = 1\]

OpenStudy (danieldjpon3):

ok?

OpenStudy (danieldjpon3):

How do we get 19?

OpenStudy (mathgeek289):

\[5+4^2+6-2^2-1\] Lets Make It Equal To 19 \[5+4^2+(6-2^2-1)\] 5 + 4^2 + (6 - 2^2 - 1) 2^2 = 4 6 - 4 - 1 = 1 5 + 4^2 + 1 4^2 = 16 5 + 16 = 19 19 + 1 = 20 So, That Cannot Be A Solution

OpenStudy (danieldjpon3):

ye man @Handsomechickenlover you prob got the question ritten down wrong

OpenStudy (handsomechickenlover):

hmmm

OpenStudy (danieldjpon3):

written*

OpenStudy (danieldjpon3):

Writen

OpenStudy (danieldjpon3):

Ugh i can't spell

OpenStudy (handsomechickenlover):

Okay

OpenStudy (handsomechickenlover):